# Hashtag Re-Appropriation for Audience Control on Recommendation-Driven Social Media Xiaohongshu (rednote)

### v1

@@ 5.4.2. Traffic Driving
Some hashtags are not used to describe the content of the post but rather serve as a means to drive traffic. We categorized these hashtags as “Attract & Irrelevant.” Since Xiaoshongshu allows up to ten hashtags for each post, these traffic-driving hashtags typically follow other hashtags that are relevant to the post content. The usage of these hashtags to drive traffic can be initiated either by the platform or by users. The most popular officially-operated hashtag on Xiaohongshu is #{CJK}UTF8gbsn笔记灵感 (Notes inspiration), which gathers a variety of popular topics in the platform. As shown in the network analysis, it is a common co-occurring hashtag with #BSF(TC) and #HG (Section 5.1.3). The hashtag was designed to increase user engagement, in return, the platform promised higher exposure for posts using this hashtag. As a result, many content creators have included this hashtag in their posts.
Another way to drive traffic is using hashtags trending in real-time on the platform. For example, P19 (man, content operator) mentioned he used #{CJK}UTF8gbsn高考 (Gaokao, National College Entrance Examination) in his meme posts to drive traffic in June, during the time of the Chinese National College Entrance Exam, which is widely discussed nationwide. This approach attracted more people to view these posts, including those who were not originally interested in the topics. Some marketing accounts also used these trending hashtags to increase views and reach a wider potential audience (Section 5.3.4).
@@ 6. Discussion
Our research uncovered the process of hashtag re-appropriation—blocking, attracting, and evolving. Given ineffective reporting mechanisms and losing trust in Xiaohongshu’s women-centered position, women communally re-appropriated hashtags to block unwanted audiences. Different user groups support this re-appropriation with varying levels. The section discusses insights from our findings: the shift towards dynamic audience control, hashtag adaptability and limitations, feminism as everyday resistance, user-driven decentralization of power, limitations and future work.
@@ 6.1. Blocking Affordance: From Static To Dynamic Audience Control
Our study highlights the ineffectiveness of Xiaohongshu’s blocking and reporting mechanisms, with many users frustrated by failed reporting attempts and the platform’s limited blocking features. In recommendation-driven platforms, many content filtering features are designed for content consumers to navigate what they want to see in their recommendation feeds, such as clicking “not interested in such content” in the recommended items, selecting interested topics in the preference setting, or providing keyword filters in the content preferences setting in TikTok  (TikTok Support, [n. d.]). However, fewer features are designed for content creators to control who sees their posts. Content creators, in particular, wanted to prevent engagement from unwanted groups while maintaining broad visibility. The need to filter unwanted audiences extends across communication platforms. For example,  Kim et al. (2021) found that, in Google Maps Place reviews, Korean speakers crafted text only understandable to fellow speakers to strategically avoid censorship and unwanted attention. Similarly, before presenting intended content, TikTok content creators employ creative strategies like starting with topics that may not be interesting for the unwanted audience to prompt them to scroll away—such as women TikTokers starting with discussing menstruation experiences to filter out male audiences  (TikTok User, [n. d.]). However, content creators’ need for robust tools of blocking unwanted audiences remains overlooked.
@@ Appendix A Interview Protocol
Here is the protocol that guides our one-hour semi-structured interview.
@@ Appendix E Post Expression Analysis
Table 5. Selected LIWC categories, abbreviations, and examples

### v2

@@ 5.4.2. Traffic Driving
Some hashtags are not used to describe the content of the post but rather serve as a means to drive traffic. We categorized these hashtags as “Attract & Irrelevant.” Since Xiaoshongshu allows up to ten hashtags for each post, these traffic-driving hashtags typically follow other hashtags that are relevant to the post content. The usage of these hashtags to drive traffic can be initiated either by the platform or by users. The most popular officially-operated hashtag on Xiaohongshu is #{CJK*}UTF8gbsn笔记灵感 (Notes inspiration), which gathers a variety of popular topics in the platform. As shown in the network analysis, it is a common co-occurring hashtag with #BSF(TC) and #HG (Section 5.1.3). The hashtag was designed to increase user engagement, in return, the platform promised higher exposure for posts using this hashtag. As a result, many content creators have included this hashtag in their posts.
Another way to drive traffic is using hashtags trending in real-time on the platform. For example, P19 (man, content operator) mentioned he used #{CJK*}UTF8gbsn高考 (Gaokao, National College Entrance Examination) in his meme posts to drive traffic in June, during the time of the Chinese National College Entrance Exam, which is widely discussed nationwide. This approach attracted more people to view these posts, including those who were not originally interested in the topics. Some marketing accounts also used these trending hashtags to increase views and reach a wider potential audience (Section 5.3.4).
@@ 6. Discussion
Our research uncovered the process of hashtag re-appropriation—blocking, attracting, and evolving. Given ineffective reporting mechanisms and losing trust in Xiaohongshu’s women-centered position, women communally re-appropriated hashtags to block unwanted audiences. Different user groups support this re-appropriation with varying levels. The section discusses insights from our findings: the shift towards dynamic audience control, hashtag adaptability and limitations, feminism as everyday resistance, user-driven decentralization of power, and limitations and future work.
@@ 6.1. Blocking Affordance: From Static To Dynamic Audience Control
Our study highlights the ineffectiveness of Xiaohongshu’s blocking and reporting mechanisms, with users frustrated by failed reporting attempts and the platform’s limited blocking features. In recommendation-driven platforms, many content filtering features are designed for content consumers to navigate what they want to see in their recommendation feeds, such as clicking “not interested in such content” in the recommended items, selecting interested topics in the preference setting, or providing keyword filters in the content preferences setting in TikTok  (TikTok Support, [n. d.]). However, fewer features are designed for content creators to control who sees their posts. Content creators, in particular, wanted to prevent engagement from unwanted groups while maintaining broad visibility. The need to filter unwanted audiences extends across communication platforms. For example,  Kim et al. (2021) found that, in Google Maps Place reviews, Korean speakers crafted text only understandable to fellow speakers to strategically avoid censorship and unwanted attention. Similarly, before presenting intended content, TikTok content creators employ creative strategies like starting with topics that may not be interesting for the unwanted audience to prompt them to scroll away—such as women TikTokers starting with discussing menstruation experiences to filter out male audiences  (TikTok User, [n. d.]). However, content creators’ need for robust tools of blocking unwanted audiences remains overlooked.
@@ Appendix A Interview Protocol
Here are the questions that guide our one-hour semi-structured interview.
@@ Appendix E Post Expression Analysis
Table 5. Selected LIWC categories, abbreviations, and examples.